# M-learner:A Flexible And Powerful Framework To Study Heterogeneous Treatment Effect In Mediation Model

### v1

@@ 4 Real Data Application
Table 4: Summary of the total treatment effect, indirect treatment effect, mediation proportion, and corresponding p-values, calculated using methods from the "mediation" package Imai et al. [2010]and the proposed approach. Med prop denotes mediation proportion:=assign:=:=indirect treatment effect/// total treatment effect. TTE denotes total treatment effect The value in parentheses is the p-value. ITE denotes indirect treatment effect. The value in parentheses is the p-value. N denotes sample size.
@@
"Mediation" NA $899$ $-0.063(0.29)$ $-0.015(0.21)$ $24.0\%(0.29)$
M-learner, no mediator Subtype1 $344$ $-0.243(0.004)$ $-0.057(0.010)$ $23.6\%(0.014)$
Subtype2 $190$ 0.003(0.950.003(0.950.003 ( 0.95) $0.005(0.77)$ $160\%(0.81)$
Subtype3 $129$ 0.10(0.42-0.10(0.42- 0.10 ( 0.42) $-0.013(0.76)$ $12.9\%(0.73)$
Subtype4 $236$ 0.145(0.030.145(0.030.145 ( 0.03) $0.009(0.47)$ $6.1\%(0.48)$
M-learner, with mediator Subtype1 $534$ $-0.15(0.016)$ $-0.030(0.010)$ $20.0\%(0.052)$
Subtype2 $365$ 0.065(0.300.065(0.300.065 ( 0.30) $0.005(0.75)$ $8.0\%(0.74)$
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Setting in no mediator model
@@
    Simple scenario in (4) , existing heterogeneity:$\eta(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$,$\kappa(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b=1$.
@@
    Complex scenario in (5), existing heterogeneity:$\eta(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$,$\kappa(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b=1$.
@@
    Simple scenario in (4), no heterogeneity,all units benefit from treatment(Global):$\eta(x)=x_{3}+x_{4}$,$\kappa(x)=1$, $b=1$.
@@
    Simple scenario in (4), no heterogeneity, no units benefit from treatment(NULL):$\eta(x)=x_{3}+x_{4}$,$\kappa(x)=0$, $b=1$.
@@
A.2 Setting in mediator model
Across all experimental designs in mediatior model, we fixed the sample size at $1,000$, with $10$ covariates generated for each unit. Subjects were randomly assigned to treatment and control groups in a $1:1$ ratio. We specified four underlying functions: $\eta_{1}(x)$ , representing the conditional mean outcome in mediator covariates model, $\eta_{2}(x)$ , representing the conditional mean outcome in covariates reponse model, and $\kappa_{1}(x)$, capturing the conditional treatment effect in mediator covariates model, and $\kappa_{2}(x)$, capturing the conditional treatment effect in covariates reponse model. All functions were defined for units under treatment ($w=1$) and control ($w=0$), for simple scenario,
@@ A.2 Setting in mediator model
    Simple scenario in (A.2), existing heterogeneity,all treatment effects via mediator:$\eta_{1}(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$,$\kappa_{1}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{1}=0$, $\eta_{2}(x)=\frac{1}{2}(x_{3}+x_{4})$,$\kappa_{2}(x)=0$, $b_{2}=1,c=1$.
@@
    Simple scenario in (A.2), existing heterogeneity, part of the treatment effect via mediator: $\eta_{1}(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$,$\kappa_{1}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{1}=0,c=1$, $\eta_{2}(x)=\frac{1}{2}(x_{3}+x_{4})$,$\kappa_{2}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{2}=1,c=1$.
@@
    Complex scenario in (A.2), existing heterogeneity, all treatment effects via mediator:$\eta_{1}(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$,$\kappa_{1}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{1}=0$, $\eta_{2}(x)=\frac{1}{2}(x_{3}+x_{4})$,$\kappa_{2}(x)=0$, $b_{2}=1,c=1$.
@@
    Complex scenario in (A.2), existing heterogeneity, part of the treatment effect via mediator: $\eta_{1}(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$,$\kappa_{1}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{1}=0,c=1$, $\eta_{2}(x)=\frac{1}{2}(x_{3}+x_{4})$,$\kappa_{2}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{2}=1,c=1$.
@@
    Simple scenario in (A.2), existing heterogeneity, $0\%$ treatment effects via mediator(NULL 1):$\eta_{1}(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$,$\kappa_{1}(x)=0$, $b_{1}=0$, $\eta_{2}(x)=\frac{1}{2}(x_{3}+x_{4})$,$\kappa_{2}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{2}=1,c=1$.
@@
    Simple scenario in (A.2), existing heterogeneity, $M$ is not mediator,All treatment effects are directly transmitted to $Y$(NULL 2):$\eta_{1}(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$,$\kappa_{1}(x)=0$, $b_{1}=0$, $\eta_{2}(x)=\frac{1}{2}(x_{3}+x_{4})$,$\kappa_{2}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{2}=1,c=0$.
@@
    Simple scenario in (A.2), no heterogeneity, all units benefit from the treatment and all treatment effects via mediator (Global):$\eta_{1}(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$,$\kappa_{1}(x)=1$, $b_{1}=0$, $\eta_{2}(x)=\frac{1}{2}(x_{3}+x_{4})$,$\kappa_{2}(x)=0$, $b_{2}=1,c=1$.
@@
A.3 Additional simulation results for M-learner
@@
Table A.7: The table summarizes the correct covariates in profiles in each scenario when sample size is $500$ in no mediators model. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$,$X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
@@
Table A.11: Comparison of true and M-learner–estimated mediation proportions and sample sizes within heterogeneous treatment effect regions across scenarios when sample size is $500$.Standard deviations are shown in parentheses. The mediation proprtion is calculated by R package "mediation". "For each set of results, the subtype exhibiting the lowest p-value for the mediation proportion is designated as the region of interest. Within this region, we estimate the mediation proportion, total treatment effect, and indirect treatment effect.
@@
Figure A.4: Comparison of $p_{leaf}$ ECDFs using RF and XGB as base learners across different scenarios with sample sizes of $1000$ and $500$.(a)results for RF without a mediator, (b) results for XGB without a mediator, (c)results for RF with a mediator (d) results for XGB with a mediator.
@@
Figure A.6: Comparison of $p_{leaf}$ ECDFs under four scenarios without a mediator, using XGB as base learner with sample sizes of $1000$ and $500$.(a)results for Simple scenario, (b) results for Complex scenario, (c)results for Global scenario (d) results for Null scenario.
@@ A.4 K-Means
Table A.12: The table summarizes the correct covariates in profiles in each scenario for K-means method in no mediators model. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$,$X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
@@
Table A.14: The table summarizes the correct covariates in profiles in each scenario for K-means method in the mediator model. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$,$X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
@@
A.5 Visualization

### v3

@@ 4 Real Data Application
Table 4: Summary of the total treatment effect, indirect treatment effect, mediation proportion, and corresponding p-values, calculated using methods from the "mediation" package Imai et al. [2010]and the proposed approach. Med prop denotes mediation proportion:=:=indirect treatment effect// total treatment effect. TTE denotes total treatment effect The value in parentheses is the p-value. ITE denotes indirect treatment effect. The value in parentheses is the p-value. N denotes sample size.
@@
"Mediation" NA $899$ 0.063(0.29)-0.063(0.29) 0.015(0.21)-0.015(0.21) $24.0\%(0.29)$
M-learner, no mediator Subtype1 $344$ 0.243(0.004)-0.243(0.004) 0.057(0.010)-0.057(0.010) $23.6\%(0.014)$
Subtype2 $190$ $0.003(0.95$) $0.005(0.77)$ $160\%(0.81)$
Subtype3 $129$ 0.10(0.42-0.10(0.42) 0.013(0.76)-0.013(0.76) $12.9\%(0.73)$
Subtype4 $236$ $0.145(0.03$) $0.009(0.47)$ $6.1\%(0.48)$
M-learner, with mediator Subtype1 $534$ 0.15(0.016)-0.15(0.016) 0.030(0.010)-0.030(0.010) $20.0\%(0.052)$
Subtype2 $365$ $0.065(0.30$) $0.005(0.75)$ $8.0\%(0.74)$
@@ 5 Discussion
Our method is unbiased when the $Y$ has linear relationship with mediator $M$, for nonlinearship, it has bias. However, with faster convergence learners, the bias of the estimator can diminish as the sample size increases.
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Assumption
Here, we introduce the assumption of the M-learner framework.
For random variables $A,B$ and $C$, let $A\perp\!\!\!\perp B|C$ denote that $A$ is conditionally independent of $B$, given $C$.
$Y(w,m)\perp\!\!\!\perp W|X$ (4)
for all $w$ and $m$.
$Y(w,m)\perp\!\!\!\perp M|X,W$ (5)
for all $w$ and $m$.
$M(w)\perp\!\!\!\perp W|X$ (6)
for all $w$.
$Y(w,m)\perp\!\!\!\perp M(w*)|X$ (7)
for all $w$,ww*,$m$.
Additionally, in the M-learner method, it assumes a randomized controlled trial (RCT) setting, the treatment assignment $W$ is independent of both the covariates $X$ and the mediator $M$, i.e.
$W\perp\!\!\!\perp X$ (8)
When there is no mediator, we can omit the $M$ in the assumptions (4) - (7).
A.2 Setting in no mediator model
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
A.3 Setting in mediator model
Across all experimental designs in mediator model, we fixed the sample size at $1,000$, with $10$ covariates generated for each unit. Subjects were randomly assigned to treatment and control groups in a $1:1$ ratio. We specified four underlying functions: $\eta_{1}(x)$ , representing the conditional mean outcome in mediator covariates model, $\eta_{2}(x)$ , representing the conditional mean outcome in covariates reponse model, and $\kappa_{1}(x)$, capturing the conditional treatment effect in mediator covariates model, and $\kappa_{2}(x)$, capturing the conditional treatment effect in covariates reponse model. All functions were defined for units under treatment ($w=1$) and control ($w=0$), for simple scenario,
@@ A.3 Setting in mediator model
[…]
@@
    Simple scenario in (A.3), existing heterogeneity, part of the treatment effect via mediator: $\eta_{1}(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$$\kappa_{1}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{1}=0,c=1$, $\eta_{2}(x)=\frac{1}{2}(x_{3}+x_{4})$$\kappa_{2}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{2}=1,c=1$.
@@
[…]
@@
    Complex scenario in (A.3), existing heterogeneity, part of the treatment effect via mediator: $\eta_{1}(x)=\frac{1}{2}(x_{1}+x_{2})+x_{3}+x_{4}$$\kappa_{1}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{1}=0,c=1$, $\eta_{2}(x)=\frac{1}{2}(x_{3}+x_{4})$$\kappa_{2}(x)=\sum_{i=1}^{2}\mathbb{I}(x_{i}>0)\cdot x_{i}$, $b_{2}=1,c=1$.
@@
[…]
@@
[…]
@@
[…]
@@
A.4 Additional simulation results for M-learner
@@
Table A.7: The table summarizes the correct covariates in profiles in each scenario when sample size is $500$ in no mediators model. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$$X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
@@
Table A.11: Comparison of true and M-learner–estimated mediation proportions and sample sizes within heterogeneous treatment effect regions across scenarios when sample size is $500$.Standard deviations are shown in parentheses. The mediation proportion is calculated by R package "mediation". "For each set of results, the subtype exhibiting the lowest p-value for the mediation proportion is designated as the region of interest. Within this region, we estimate the mediation proportion, total treatment effect, and indirect treatment effect.
@@
Figure A.4: Threshold location distribution plot for Simple-All scenario. dot in the figure represents the threshold of heterogeneous region for 100 replications, red line represents the density line. If a variable was not selected or the threshold exceeded 1, we assigned a value of 1. (a) Simple-All, (b) Simple-Part, (c) Complex-All, (d) Complex-Part.
Figure A.5: Comparison of $p_{leaf}$ ECDFs using RF and XGB as base learners across different scenarios with sample sizes of $1000$ and 500500.(a)results for RF without a mediator, (b) results for XGB without a mediator, (c)results for RF with a mediator (d) results for XGB with a mediator.
@@
Figure A.7: Comparison of $p_{leaf}$ ECDFs under four scenarios without a mediator, using XGB as base learner with sample sizes of $1000$ and 500500.(a)results for Simple scenario, (b) results for Complex scenario, (c)results for Global scenario (d) results for Null scenario.
@@ A.5 K-Means
Table A.12: The table summarizes the correct covariates in profiles in each scenario for K-means method in no mediators model. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$$X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
@@
Table A.14: The table summarizes the correct covariates in profiles in each scenario for K-means method in the mediator model. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$$X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
@@
A.6 Sensitivity analysis
A.6.1 Noise
To evaluate the robustness of the proposed method under varying levels of noise, we introduced additive Gaussian noise with different variances. The noise levels were categorized as follows:
  • low noise (original setting): $N(0,0.01)$;
  • moderate noise: $N(0,0.1)$;
  • high noise: $N(0,1)$ .
All other experimental settings were kept identical to Simple-All and Complex-All in A.10.2. For both the moderate and high noise settings, calibration is performed using the Simple-Null2 distribution obtained under the low noise scenario (To simulate real-world data characteristics, we perform calibration using data with a low level of noise). What’s more, we want to evaluate the error of the selected heterogeneous region. Here, we define the threshold error, for each run, we extracted the decision threshold defining the heterogeneous region and computed the average of the squared threshold values. If a variable was not selected or the threshold exceeded 1, we assigned a value of 1.
From Table A.17, A.21,A.19 and Figure A.11, suggest that noise level does have some effect on the model’s performance in the Simple-All scenario. In particular, performance tends to degrade slightly under higher levels of noise.
From Table A.18, A.22,A.20 and Figure A.12, suggest that noise level does have some effect on the model’s performance in the Complex-All scenario. Similar with Simple-All scenario, performance tends to degrade slightly under higher levels of noise.
These experimental findings provide insights into the effect of noise on model performance. While noise introduces some variability, its overall impact is limited. Notably, under high-noise conditions, the learner based on XGBoost even outperforms the Random Forest learner under low-noise settings.
Table A.17: The table summarizes the correct covariates in profiles in scenario Simple-All for different noise levels with XGB learner. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$$X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
Covariates $X_{1}$ $X_{2}$ $X_{1}\&X_{2}$
Low 100 100 100
Moderate 96 98 95
High 91 94 85
Table A.18: The table summarizes the correct covariates in profiles in scenario Complex-All for different noise levels with XGB learner. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$$X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
Covariates $X_{1}$ $X_{2}$ $X_{1}\&X_{2}$
Low 98 99 97
Moderate 100 100 100
High 83 87 72
Table A.19: The table summarizes the boundary osf selected heterogeneous regions in profiles in scenario Simple-All for different noise levels with XGB learner. Values are the mean threshold errors over 100 replications, with standard deviations shown in parentheses.
Covariates $X_{1}$ $X_{2}$
Low 0.056(0.194) 0.059(0.162)
Moderate 0.043(0.173) 0.047(0.153)
High 0.111(0.285) 0.106(0.289)
Table A.20: The table summarizes the boundary osf selected heterogeneous regions in profiles in scenario Complex-All for different noise levels with XGB learner. Values are the mean threshold errors over 100 replications, with standard deviations shown in parentheses.
Covariates $X_{1}$ $X_{2}$
Low 0.425(0.489) 0.475(0.481)
Moderate 0.372(0.483) 0.531(0.529)
High 0.419(0.478) 0.561(0.475)
Table A.21: The table summarizes the distribution of the number of variables included in the selected decision trees across 100 simulation runs with different noises levels in Simple-All scenario. Each value represents the frequency with which a specific number of variables was selected over the course of the experiments. Selecting zero covariates in a given experimental run indicates that the decision tree’s classification was not supported by the calibration process and was therefore rejected.
Number of covariates 0 1 2 3 4 $\geq$5
Low 0 0 58 33 9 0
Moderate 1 2 79 15 3 0
High 0 12 63 18 7 0
Table A.22: The table summarizes the distribution of the number of variables included in the selected decision trees across 100 simulation runs with different noises levels in Complex-All scenario. Each value represents the frequency with which a specific number of variables was selected over the course of the experiments. Selecting zero covariates in a given experimental run indicates that the decision tree’s classification was not supported by the calibration process and was therefore rejected.
Number of covariates 0 1 2 3 4 $\geq$5
Low 0 0 58 33 9 0
Moderate 0 0 19 47 34 0
High 2 6 33 40 16 3
Figure A.11: Empirical cumulative distribution functions (ECDF) of $p_{leaf}$ under different noises levels in Simple-All scenario. These results illustrate the sensitivity of each base learner to underlying treatment effect heterogeneity across varying levels of complexity.
Figure A.12: Empirical cumulative distribution functions (ECDF) of $p_{leaf}$ under different noises levels in Complex-All scenario. These results illustrate the sensitivity of each base learner to underlying treatment effect heterogeneity across varying levels of complexity.
Figure A.13: Threshold location distribution plot for Simple-All scenario. dot in the figure represents the threshold of heterogeneous region for 100 replications, red line represents the density line. If a variable was not selected or the threshold exceeded 1, we assigned a value of 1. (a) Low noise, (b) Moderate noise, (c) High noise.
Figure A.14: Threshold location distribution plot for Complex-All scenario. dot in the figure represents the threshold of heterogeneous region for 100 replications, red line represents the density line. If a variable was not selected or the threshold exceeded 1, we assigned a value of 1. (a) Low noise, (b) Moderate noise, (c) High noise.
A.6.2 dimension of projection
Here, we evaluate the robustness of the dimension of projection. t-SNE only support reduce dimension to $2$ or $3$. Here we compare the results in Simple-All and Complex-All scenario with different dimension of projections. For UMAP, it support reduce to higher dimension, we compare the results with $2,5$ and $10$ in Simple-All and Complex-All scenario.
As shown in the Table A.23, the performance of t-SNE appears insensitive to the choice of projection dimension, while UMAP exhibits a clear dependence on it. In particular, UMAP achieves optimal performance when the projection dimension is set to 5, outperforming both the 2- and 10-dimensional settings. Furthermore, even the best t-SNE result is consistently outperformed by UMAP under its optimal configuration.
Table A.23: The table summarizes the correct covariates in profiles in each scenario. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$, $X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
Simple-All Complex-All
Covariates $X_{1}$ $X_{2}$ $X_{1}\&X_{2}$ $X_{1}$ $X_{2}$ $X_{1}\&X_{2}$
t-SNE(2) 100 100 100 100 100 100
t-SNE(3) 100 100 100 100 100 100
UMAP(2) 97 98 96 91 86 82
UMAP(5) 99 100 99 96 92 91
UMAP(10) 94 94 89 89 84 79
A.6.3 Number of maximum clusters
To investigate the impact of the maximum number of clusters on the results, we evaluate the effect of setting the maximum cluster number to 2, 5, and 10 in both the Simple-All and Complex-All scenarios. From Table A.24 and A.25, we observe that the maximum number of clusters $K$ influences the results to some extent, but this effect becomes negligible once $K$ exceeds a certain threshold. Therefore, in practice, we recommend choosing an appropriate value of $K$ based on computational efficiency. To ensure optimal performance, we suggest selecting the largest feasible $K$ within resource constraints.
Table A.24: The table summarizes the correct covariates in profiles in Simple-All and Complex-All scenario for different maximum clusters $K$. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$, $X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
Simple-All Complex-All
Covariates $X_{1}$ $X_{2}$ $X_{1}\&X_{2}$ $X_{1}$ $X_{2}$ $X_{1}\&X_{2}$
$K=2$ 87 83 75 81 87 72
$K=5$ 100 100 100 98 99 97
$K=10$ 100 100 100 100 100 100
Base Learner Simple-All Complex-All
Number of covariates 0 1 2 3 4 $\geq$5 0 1 2 3 4 $\geq$5
$K=2$ 5 12 66 14 3 0 4 12 54 21 8 1
$K=5$ 3 1 46 39 11 0 0 1 23 43 33 0
$K=10$ 37 6 16 23 13 0 0 0 9 38 53 0
A.7 Comparison of projection methods
The core idea of our M-learner algorithm is that individuals with similar treatment effects should be considered more similar. Based on this intuition, we construct a pairwise treatment effect distance matrix that reflects heterogeneity between any two individuals. We then use t-SNE to project this matrix into a Euclidean space for clustering.
t-SNE (similarly UMAP) is particularly suitable here because it is designed to preserve local pairwise similarities during projection, which aligns with our goal of preserving treatment effect similarity. Unlike PCA, which performs linear projections and assumes Euclidean structure in the original space, our distance matrix is derived from treatment effect heterogeneity—not raw features—making non-linear methods like t-SNE or UMAP more appropriate. Notably, in our study, t-SNE directly operates on pairwise distances, which further justifies its application. In our study, t-SNE is employed to project the treatment effect distance matrix onto a low-dimensional Euclidean space while preserving local neighborhood structures. This approach facilitates the effective identification of potential subgroups, which are subsequently extracted through clustering.
Moreover, we also conducted sensitivity analyses using alternative projection methods such as UMAP(the goal in this step is not to reduce dimensionality per se, but rather to project the treatment distance matrix into a space that preserves their relative distances for downstream clustering. PCA focuses on preserving global variance rather than local or relational structure, it is not suitable for this purpose), and the results remained qualitatively consistent.
In this section, we compare the results for two methods. In this experiment, we adopted XGBoost as the sole base learner. The configurations for both the without mediator and with mediator scenarios followed those of Experiment A.10.1 and A.10.2, with the only difference being the choice of projection method.
From Table A.26, A.27 and Figure A.15 results, we can think there is little difference between two projection methods when there is no mediators. However, from the results from Table A.28, A.29 and Figure A.15 when there exists a mediator, UMAP perform worse than t-SNE in complex scenarios. So we choose t-SNE as the projection method.
Table A.26: The table summarizes the correct covariates in profiles in each scenario. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$$X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
Projection method UMAP t-SNE
Covariates $X_{1}$ $X_{2}$ $X_{1}\&X_{2}$ $X_{1}$ $X_{2}$ $X_{1}\&X_{2}$
Simple 100 100 100 100 100 100
Complex 72 74 68 70 70 68
Global 0 1 0 0 2 0
Null 2 2 1 1 4 1
Table A.27: The table summarizes the distribution of the number of variables included in the selected decision trees across 100 simulation runs when there is no mediator and sample size is $1000$. Each value represents the frequency with which a specific number of variables was selected over the course of the experiments. The base learner is XGBoost. This table compares UMAP and t-SNE as two distinct projection methods under four distinct experimental scenarios Selecting zero covariates in a given experimental run indicates that the decision tree’s classification was not supported by the calibration process and was therefore rejected.
Projection method UMAP t-SNE
Number of covariates 0 1 2 3 $\geq$4 0 1 2 3 $\geq$4
Simple 0 0 62 28 10 0 0 57 28 15
Complex 21 1 15 29 14 28 1 7 30 34
Global 95 1 3 1 0 98 0 2 0 0
Null 90 2 4 4 0 90 3 3 3 1
Table A.28: The table summarizes the correct covariates in profiles in each scenario. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$, $X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
UMAP t-SNE
Covariates $X_{1}$ $X_{2}$ $X_{1}\&X_{2}$ $X_{1}$ $X_{2}$ $X_{1}\&X_{2}$
Simple-All 97 98 96 100 100 100
Simple-Part 100 99 99 100 100 100
Complex-All 91 86 82 98 99 97
Complex-Part 79 69 61 95 97 93
Simple-Null1 6 2 2 8 4 2
Simple-Null2 2 3 0 2 1 0
Simple-Global 1 3 1 3 5 1
Table A.29: The table summarizes the distribution of the number of variables included in the selected decision trees across 100 simulation runs. Each value represents the frequency with which a specific number of variables was selected over the course of the experiments. The base learner is XGBoost. This table compares UMAP and t-SNE as two distinct projection methods under seven distinct experimental scenarios. Selecting zero covariates in a given experimental run indicates that the decision tree’s classification was not supported by the calibration process and was therefore rejected.
Projection method UMAP t-SNE
Number of covariates 0 1 2 3 4 $\geq$5 0 1 2 3 4 $\geq$5
Simple-All 1 3 61 17 15 3 0 0 58 33 9 0
Simple-Part 0 1 68 23 8 0 0 0 54 40 6 0
Complex-All 5 6 25 42 22 0 0 1 23 43 33 0
Complex-Part 13 10 25 34 18 0 1 1 14 49 35 0
Simple-Null1 86 6 5 1 1 1 83 1 8 6 1 1
Simple-Null2 91 2 4 2 1 0 90 5 3 2 0 0
Simple-Global 90 2 7 1 0 0 87 1 6 4 2 0
Figure A.15: Empirical cumulative distribution functions (ECDF) of $p_{leaf}$ for projection comparison.(a)results for UMAP without a mediator, (b) results for UMAP with a mediator.
A.8 Different clustering methods
We compare different clustering methods, such as K-Medoids and DBSCAN in Simple scenario (no mediator setting). DBSCAN does not depend on predefined numbers of clusters, we set eps is 1, minPts is 20 with R dbscan function. For K-Medoids, similar to K-Means, we preset the number of clusters to range from $2$ to $5$. Here, we compare the performance in simple scenario in non-mediated setting. from A.16, K-Medoids method close to K-Means, and DBSCAN is not acceptable.
Figure A.16: Empirical cumulative distribution functions (ECDF) of $p_{leaf}$ for clustering comparison.(a)results for K-Medoids, (b) results for DBSCAN. Cov1 represnets "X1", Cov2 represnets "X2".
A.9 Comparison in non-mediators setting
When no mediators are present, our method is equivalent to the T-learner in the estimation of the CATTE. In this subsection, we compare the performance of several popular methods for subgroup identification (using the same clustering and balancing approach described later in this paper), including X-learner, R-learner, and a neural network-based method, TARNet. To ensure a fair comparison, all learner-based methods were implemented using XGBoost and estimated via the econml Python library. The TARNet model was implemented using the DragonNet architecture from the causalml library, with the propensity score weighting component removed.
To evaluate subgroup identification, we visualized the estimated thresholds defining heterogeneous subgroups across methods. Given that the true subgroup in our simulation is characterized by $X_{1}>0$ and $X_{2}>0$, we extracted the estimated boundaries (i.e., a, b such that the identified subgroup is $X_{1}>a,X_{2}>b$). For cases where a method did not select a covariate or the estimated threshold exceeded 1, we set the corresponding value to 1. The results are shown in Figures A.17 - A.23
Figure A.17: Visualization of heterogeneous subgroup thresholds estimated by the M-learner. Each point represents the threshold identified in a single simulation run. Thresholds greater than 1 or cases where the corresponding variable was not selected are recorded as 1. The red line indicates the estimated density. Cov1 represnets "X1", Cov2 represnets "X2".
Figure A.18: Visualization of heterogeneous subgroup thresholds estimated by the X-learner. Each point represents the threshold identified in a single simulation run. Thresholds greater than 1 or cases where the corresponding variable was not selected are recorded as 1. The red line indicates the estimated density.
Figure A.19: Visualization of heterogeneous subgroup thresholds estimated by the R-learner. Each point represents the threshold identified in a single simulation run. Thresholds greater than 1 or cases where the corresponding variable was not selected are recorded as 1. The red line indicates the estimated density.
Figure A.20: Visualization of heterogeneous subgroup thresholds estimated by the TARNet. Each point represents the threshold identified in a single simulation run. Thresholds greater than 1 or cases where the corresponding variable was not selected are recorded as 1. The red line indicates the estimated density.
Figure A.21: Empirical cumulative distribution functions (ECDF) of $p_{leaf}$ for X-learner.
Figure A.22: Empirical cumulative distribution functions (ECDF) of $p_{leaf}$ for R-learner.
Figure A.23: Empirical cumulative distribution functions (ECDF) of $p_{leaf}$ for TARNet.
A.10 Visualization